2x-5y=7
8=x-4y
I have to solve using substitution. My answer is (20,3)which means x=20 and y=3.

Found 2 solutions by Alan3354, richwmiller:
Answer by Alan3354(69443)

2x-5y=7
8=x-4y --> x = 4y + 8
Sub for x into 1st eqn:
2(4y+8) - 5y = 7
8y+16 - 5y = 7
3y = -9
y = -3
-------
x = -4
------
I have to solve using substitution. My answer is (20,3)which means x=20 and y=3.
Check (20,3)
2x-5y=7
2*20 - 5*3 =? 7
It doesn't.
-------------
It's (-4,-3)

Answer by richwmiller(17219)

It's (-4,-3 )
2x-5y=7
x=4(y+2)
2*4(y+2)-5y=7
8y+16-5y=7
3y=-9
y=-3
x=4(y+2)
x=4(-3+2)
x=-4
(-4,-3)